OpenStudy (anonymous):

What are the exact solutions of x2 = 4 − 7x?

OpenStudy (anonymous):

is x2 really x^2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x^2+7x=4 8x^2=4 x^2= 4/8 x^2= 1/2 x=\[\sqrt{1/2}\]

OpenStudy (anonymous):

thats none of the answers lol

OpenStudy (anonymous):

huh then i got wrong maybe dpalnc can help

OpenStudy (anonymous):

get the quadratic in standard form first: x^2 + 7x - 4 = 0 a=1 b=7 c=-4 use quadratic formula: \[\large x=\frac{-b\pm\sqrt{b^2-4ac}}{2a} \]

OpenStudy (anonymous):

i see that but its so confusing...

OpenStudy (anonymous):

\[\large x =\frac{-b\pm \sqrt{b^2-4ac}}{2a}=\frac{-(7)\pm \sqrt{7^2-4(1)(-4)}}{2(1)} \]